A railling of 288 meter is required for fencing a semicircular park.find the area of the park
Answer by swincher4391(1107)

288 = pi*r + 2*r
288 = (2+pi) * r
r = 56.01
A = pi*r^2 = pi * 56.01^2 = 9856m^2